Please find a counterexample to the statement: The square of a number added to the sum of twice the number and 8 is a prime number. Justify your solution by supporting it with step by step work.
Thank you very much

Answer by themathtutor2009(81)

Counterexample: x=2


x^2+2x+8 = 2^2+2(2)+8=4+4+8=16 is not prime